OpenStudy (anonymous):

find indefinite integral of (x^2)*(sqrt(5x+4))

hartnn (hartnn):

hint : substitute 5x+4 = u.

OpenStudy (anonymous):

\[x ^{2}\times \sqrt{5x+4}\]

OpenStudy (anonymous):

I did that but it got really messy really fast

hartnn (hartnn):

x= (u-4) /5 so, x^2 =... ? also dx= du/5

OpenStudy (anonymous):